# Hidden yet quantifiable: A lower bound for confounding strength using randomized trials

### v3

@@ A.2 Nested design
In a nested trial design, the randomized trial is embedded in a cohort of eligible people who are proposed to participate in the trial, but if they refuse, they are still included in the observational study. Two concrete examples of nested designs are the Women Health Initiative [1] and the recent study on Medicaid [14].
@@
where the sampling probability $\mathbb{Q}(S\mid X)$ can be identified under a nested study design [11].
@@ A.3 Limitations of MSM
We discuss here the intuition behind the tightness of $\hat{\Gamma}_{\texttt{LB}}$ in the infinite-sample limit, though it carries over to finite samples.
Without loss of generality, we focus on the lower bound derived from $\hat{\phi}_{\widetilde{\mathrm{os}}}$ and assume that the unobserved confounding biases the average treatment effect upwards, i.e.
$\mu^{-}_{\Gamma=1}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu^{+}_{\Gamma=1}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})\geq\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{\widetilde{os}}}_{\operatorname{full}}),$
where $\mu^{-}_{\Gamma=1}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})$ and $\mu^{+}_{\Gamma=1}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})$ are the IPW estimates of ATE on the (restricted) marginal distribution of the observational study. Further, we define the infinite-sample lower bound as
@@
which is, in words, the value of $\Gamma$ such that the sensitivity bounds include the true ATE. Formally, it holds that
$\Gamma_{\texttt{LB}}$ $=\inf\left\{\Gamma:\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})\in[\mu_{\Gamma}^{+}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}),\mu_{\Gamma}^{-}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})]\right\}$
$=\inf\left\{\Gamma:\mu_{\Gamma}^{+}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})\quad\text{or}\quad\mu_{\Gamma}^{-}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})\right\}$
$=\inf\left\{\Gamma:\mu_{\Gamma}^{-}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})\right\},$
where the second equality follows from the monotonicity of the sensitivity bounds and the last equality from the assumption of the bias direction. Since the sensitivity bounds are continuous and strictly increasing, the set is non-empty and contains one element.
@@
  • If $\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}=\underset{\tilde{\mathbb{P}}_{\operatorname{full}}\in\mathcal{E}(\mathbb{P}^{\widetilde{\mathrm{os}}},\Gamma^{\star})}{\text{argmax}}\,\mu(X,\tilde{\mathbb{P}}_{\operatorname{full}})$, it holds that $\mu_{\Gamma^{\star}}^{+}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})$ and $\Delta=0$. This case is achieved when $U=(Y(1),Y(0))$ (see Dorn and Guo [17] for a closed-form solution of the sensitivity bounds). Intuitively, the MSM does not place any assumptions on the form of the confounder, and the worst-case $\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}$ is achieved when the confounder is equal to the potential outcomes.
  • If $U\perp\!\!\!\perp(Y(1),Y(0))$, it holds that $\mu^{-}_{\Gamma=1}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu^{+}_{\Gamma=1}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})=\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{\widetilde{os}}}_{\operatorname{full}})$. Hence, $\Delta=\Gamma^{\star}-1$ and the lower bound can be arbitrarily loose.
@@ B.1 Synthetic experiments
are respectively the lower and upper bounds on the full propensity score under the MSM. By choosing $t(X)=\frac{e(X)-\ell(X)}{u(X)-\ell(X)}$ in our data-generating process in Section 4.1 where $U\sim U(0,1)$, we ensure that $\mathbb{E}_{\mathbb{P}^{\operatorname{os}}}[e^{+}(X,U)\mid X]=e(X)$. We note that this is different from [35, 32, 47] where they choose a fixed threshold $t(X)=1/2$, resulting in a data distribution that does not satisfy the MSM. For all synthetic experiments, we estimate the propensity score using logistic regression. Further, we set $n_{\text{bootstrap}}=100,\sigma^{2}_{Y}=0.1,\alpha=0.05$, and report the mean and standard error over 20 runs.
For the test $\hat{\phi}_{\widetilde{\mathrm{os}}}$, we use the sensitivity bound estimator QB [17], and fit the quantile function using quantile forest regression [41]. For the test $\hat{\phi}_{\mathrm{rct}}$, we use the sensitivity bound estimator B-Learner [47]. We fit the quantile function using quantile forest regression [41], and the outcome model using a random forest regressor.
@@ B.2.1 Subsampling procedure
Algorithm 2 Randomized Trial Rejection Sampling [37]
@@
7:         if $K_{i}>\frac{\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=t_{i}\mid U_{i})}{\hat{\mathbb{P}^{\operatorname{rct}}}(T=t_{i})M}$ then
@@
where we estimate $\hat{\pi}=\hat{\mathbb{P}^{\operatorname{rct}}}(T=1)$ from $D\setminus D_{\operatorname{rct}}$. For a continuous confounder $U$ positively correlated with $Y(1)$, we use the following full propensity score in Algorithm 2:
$\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1\mid U)=\begin{cases}\ell&\text{if }U>Q_{\hat{q}^{*}}(U),\\ u&\text{if }U<Q_{\hat{q}^{*}}(U),\end{cases}$ (9)
where $Q_{q}(U)=\inf\left\{z\in\mathbb{R}:\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(U\leq z)\geq q\right\}$ is the $q$-th quantile of the marginal distribution of $U$. Since the propensity score does not depend on $X$, using $\hat{q}^{\star}=\frac{u-\hat{\pi}}{u-\ell}$ makes sure that
$e(X)=\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1|X)=\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1)=\mathbb{E}_{U}\left[\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1|U)\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(U)\right]=\hat{\mathbb{P}^{\operatorname{rct}}}(T=1)$
and hence the subsampled dataset, $D_{\mathrm{os}}$, satisfies $\mathcal{E}(\mathbb{P}^{\operatorname{os}},\Gamma^{\star})$. Note that this is equivalent to enforcing the same marginal propensity score before and after the subsampling. For a negatively correlated continuous confounder, we choose $\hat{q}^{\star}=\frac{\hat{\pi}-\ell}{u-\ell}$ and change the direction of the inequalities in Equation 9.
Further, for a positively correlated binary confounder, we use the following full propensity score in Algorithm 2:
@@
$\hat{\mathbb{P}^{\operatorname{rct}}}(U=1)$ $=\frac{u-\hat{\pi}}{u-\ell},$
and then applying Algorithm 2 with the full propensity score from Equation 10, we again obtain $e(X)=\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1)=\hat{\mathbb{P}^{\operatorname{rct}}}(T=1)$. For a negatively correlated binary confounder, we first enforce $\hat{\mathbb{P}^{\operatorname{rct}}}(U=1)=\frac{\hat{\pi}-\ell}{u-\ell}$ and swap $\ell$ for $u$ in Equation 10, and vice versa.
The subsampling procedure in Algorithm 2 allows for the construction of observational datasets where the causal effect is identifiable non-parametrically, in contrast to previous approaches, such as [21], that do not guarantee identifiability after subsampling. In Algorithm 2, we note that $M=\frac{\max_{i\in\{1,\ldots,n\}}\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}\left(T=t_{i}\mid U_{i}\right)}{\min_{i\in\{1,\ldots,n\}}\hat{\mathbb{P}^{\operatorname{rct}}}\left(T=t_{i}\right)}$ in the limit satisfies $M\geq\sup_{T,U}\frac{\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T\mid U)}{\hat{\mathbb{P}^{\operatorname{rct}}}(T)}$, which is required to ensure that the causal effect is identifiable in the subsampled dataset (see Theorem 3.2 in [37]). We empirically observe that the subsampling procedure approximately discards half of the instances from $D\setminus D_{\operatorname{rct}}$.
@@ B.2.2 Datasets details
  • Hillstrom’s MineThatData Email data [27]. The Hillstrom dataset contains records of 64,000 customers who purchased within the last twelve months. They were part of an e-mail campaign to assess the effectiveness of distinct campaign strategies. Two treatment groups, “Men’s” and “Women’s” email campaigns, and a control group were established. Treatments were randomly assigned. Our analysis primarily focuses on a combined treatment group, which constitutes roughly 66% of the dataset. While the original dataset has different outcomes, we looked at the dollars spent in the two weeks post-campaign. The dataset provides data on recent purchase patterns (Recency), annual spending categories (History Segment) and values (History), merchandise type, either Mens (Mens) or Womens (Womens), geographical location via zip code (Zip Code), newcomer status (Newbie), and purchasing avenues (Channel). After subsampling, we end up with a randomized trial of size $n_{\operatorname{rct}}=7680$ and an observational dataset of size $n_{\operatorname{os}}=20500$. 2.3 is enforced by excluding urban zip codes in the trial.
@@
  • VOTE dataset [22]. The VOTE dataset studies the effect of social pressure on voting behaviors among Michigan’s registered voters, focusing on those who voted in the prior election and met certain criteria. The primary outcome is a binary variable indicating whether the letter recipients voted. In this randomized trial, participants were allocated to a control group or one of four treatment groups. The treatment groups received distinct letters, each varying in social pressure intensity, aimed at encouraging voting. The most persuasive letter provided insight into the recipient’s neighbors’ voting patterns from the previous two elections and implied updates on neighbors’ subsequent voting actions in future letters. Using the split in [55], we incorporated roughly 190,000 samples in the control group, and we kept the treatment group with the strongest letter, leaving about 38,000 samples. We retained preprocessed features like age, household size, gender, and two scores reflecting past voting habits and the voting patterns of neighbors. After subsampling, we end up with a randomized trial of size $n_{\operatorname{rct}}=10650$ and an observational dataset of size $n_{\operatorname{os}}=36800$. We discard households with more than 4 participants to enforce 2.3.
@@
  • Tennessee STAR Project [63]. The Tennessee STAR experiment, initiated in 1985, was a randomized study examining the impact of class size on students’ standardized test scores, tracking them from kindergarten through third grade. Initially, students and teachers were randomly placed into class sizes, intending to maintain these conditions throughout the study. We follow the dataset preprocessing outlined in [34]. Their analysis concentrates on two conditions: small classes (13-17 students) and regular-sized classes (22-25 students). They used the class size at first grade as the treatment variable, observing 4,509 students. Their outcome aggregates scores from listening, reading, and math tests at the end of the first grade. After excluding students with missing values, the final sample consisted of 4,218 students: 1,805 in small classes (treatment) and 2,413 in regular-sized classes (control). The observed features for each student are gender, race, birth month, birthday, birth year, free lunch given or not and teacher ID. After subsampling, we end up with a randomized trial of size $n_{\operatorname{rct}}=600$ and an observational dataset of size $n_{\operatorname{os}}=1800$. For the STAR project, we keep inner-city and suburban students but remove urban and rural ones to enforce 2.3.
@@
We use QB [17] for the test $\hat{\phi}_{\widetilde{\mathrm{os}}}$. For continuous outcomes (Hillstrom and STAR), we fit a random forest regressor for the quantile functions, while we leverage the closed-form solution for the quantiles in the binary case (VOTE). For the test $\hat{\phi}_{\mathrm{rct}}$ we use the B-Learner [47], fitting respective random forest regressors for the quantiles, the outcome and the bounds models. For the binary case, we use the closed-form solution for the quantiles and fit the outcome and bounds models with XGBoost [4]. We always train a logistic regressor for the propensity score. We report mean and standard error over 15 runs and set $n_{\text{bootstrap}}=200,\alpha=0.05$ for all experiments.
@@ B.3 Women’s Health Initiative
We binarize a composite outcome, called the “global index”, in our analysis, where Y = 1 if coronary heart disease or stroke was observed in the first seven years of follow-up, and Y = 0 otherwise. Note that Y = 0 could also occur from censoring. To establish treatment and control groups in the observational study, we use questionnaire data in which participants confirm or deny usage of combination hormones (i.e. both estrogen and progesterone) in the first three years. Using this procedure, we end up with a total of $n_{\operatorname{os}}=33,511$ patients. Finally, we restrict the set of covariates used to those that are measured in both the RCT and the observational study. In particular, we use as covariates only those measured in both the RCT and observational study, and we further restrict them to those identified as significant in epidemiological literature, such as in [49]. Specifically, the covariates in our analysis are: AGE, ETHNIC_White, BMI, SMOKING_Past_Smoker, SMOKING_Current_Smoker, EDUC_x_College_graduate_or_Baccalaureate Degree, EDUC_x_Some_post-graduate_or_professional, MENO, PHYSFUN. The data used is available on BIOLINCC.
@@
We train logistic regression for both outcome models and propensity score. We use as sensitivity bounds DVDS [18] for the test $\hat{\phi}_{\widetilde{\mathrm{os}}}$, and B-Learner for the test $\hat{\phi}_{\mathrm{rct}}$. We test for confounding in one direction, i.e. we only compute the test statistic $\hat{T}_{\Gamma}^{+}$. We set $n_{\text{bootstrap}}=100$ and $\alpha=0.05$ for all experiments.

### v4

@@ A.2 Nested design
In a nested trial design, the randomized trial is embedded in a cohort of eligible people who are proposed to participate in the trial, but if they refuse, they are still included in the observational study. Two concrete examples of nested designs are the Women Health Initiative [anderson2003implementation] and the recent study on Medicaid [degtiar2021conditional].
@@
where the sampling probability $\mathbb{Q}(S\mid X)$ can be identified under a nested study design [dahabreh2021study].
@@ A.3 Limitations of MSM
We discuss here the intuition behind the tightness of $\hat{\Gamma}_{\texttt{LB}}$ in the infinite-sample limit, though it carries over to finite samples. Without loss of generality, we focus on the lower bound derived from $\hat{\phi}_{\widetilde{\mathrm{os}}}$ and assume that the unobserved confounding biases the average treatment effect upwards, i.e.
[…]
@@
[…]
@@
Case 1.
If
$\mathbb{P}^{\operatorname{\widetilde{os}}}_{\operatorname{full}}\in\underset{\tilde{\mathbb{P}}_{\operatorname{full}}\in\mathcal{E}(\mathbb{P}^{\widetilde{\mathrm{os}}},\Gamma^{\star})}{\operatorname{argmin}}\,\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\tilde{\mathbb{P}}_{\operatorname{full}}),$
[…]
Case 2.
[…]
@@ B.1 Synthetic experiments
are respectively the lower and upper bounds on the full propensity score under the MSM. By choosing $t(X)=\frac{e(X)-\ell(X)}{u(X)-\ell(X)}$ in our data-generating process in Section 4.1 where $U\sim U(0,1)$, we ensure that $\mathbb{E}_{\mathbb{P}^{\operatorname{os}}}[e^{+}(X,U)\mid X]=e(X)$. We note that this is different from [kallus2019interval, jesson2021quantifying, oprescu2023b] where they choose a fixed threshold $t(X)=1/2$, resulting in a data distribution that does not satisfy the MSM. For all synthetic experiments, we estimate the propensity score using logistic regression. Further, we set $n_{\text{bootstrap}}=100,\sigma^{2}_{Y}=0.1,\alpha=0.05$, and report the mean and standard error over 20 runs.
For the test $\hat{\phi}_{\widetilde{\mathrm{os}}}$, we use the sensitivity bound estimator QB [dorn2022sharp], and fit the quantile function using quantile forest regression [meinshausen2006quantile]. For the test $\hat{\phi}_{\mathrm{rct}}$, we use the sensitivity bound estimator B-Learner [oprescu2023b]. We fit the quantile function using quantile forest regression [meinshausen2006quantile], and the outcome model using a random forest regressor.
@@ B.2.1 Subsampling procedure
Algorithm 2 Randomized Trial Rejection Sampling [keith2023rct]
@@
7:   if $K_{i}>\frac{\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=T_{i}\mid U_{i})}{\widehat{\mathbb{P}}(T=t_{i})M}$ then
@@
where we estimate $\hat{\pi}=\widehat{\mathbb{P}}(T=1)$ from the split $D\setminus D_{\operatorname{rct}}$ before subsampling. For a continuous confounder $U$ positively correlated with $Y(1)$, we use the following full propensity score in Algorithm 2:
$\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1\mid U)=\begin{cases}\ell&\text{if }U>Q_{\hat{q}^{*}}(U),\\ u&\text{if }U\leq Q_{\hat{q}^{*}}(U),\end{cases}$ (9)
where $Q_{q}(U)=\inf\left\{z\in\mathbb{R}:{\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(U\leq z)}\geq q\right\}$ is the $q$-th quantile of the marginal distribution of $U$ in $D\setminus D_{\operatorname{rct}}$ before subsampling. Since the propensity score does not depend on $X$, using $\hat{q}^{\star}=\frac{\hat{\pi}-\ell}{u-\ell}$ makes sure that
$e(X)=\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1|X)=\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1)=\hat{\pi}$
and hence the target full propensity score satisfies the marginal constraint. Note that this is equivalent to enforcing the same marginal propensity score before and after the subsampling. For a negatively correlated continuous confounder, we choose $\hat{q}^{\star}=\frac{u-\hat{\pi}}{u-\ell}$ and change the direction of the inequalities in Equation 9.
Further, for a positively correlated binary confounder, we use the following propensity score in Algorithm 2:
@@
$\widehat{\mathbb{P}}(U=1)$ $=\frac{u-\hat{\pi}}{u-\ell},$
and then applying Algorithm 2 with the propensity score from Equation 10, we again obtain $e(X)=\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1)=\hat{\pi}$. For a negatively correlated binary confounder, we first enforce $\widehat{\mathbb{P}}(U=1)=\frac{\hat{\pi}-\ell}{u-\ell}$ and swap $\ell$ for $u$ in Equation 10, and vice versa.
The subsampling procedure in Algorithm 2 allows for the construction of observational datasets where the causal effect is identifiable non-parametrically, in contrast to previous approaches, such as [gentzel2021and], that do not guarantee identifiability after subsampling. In Algorithm 2, we note that $M=\frac{\max_{i\in\{1,\ldots,n\}}\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}\left(T=t_{i}\mid U_{i}\right)}{\min_{i\in\{1,\ldots,n\}}\widehat{\mathbb{P}}\left(T=t_{i}\right)}$, which is required to ensure that the causal effect is identifiable in the subsampled dataset (see Theorem 3.2 in [keith2023rct]). We empirically observe that the subsampling procedure discards half of the instances from $D\setminus D_{\operatorname{rct}}$.
@@ B.2.2 Datasets details
  • Hillstrom’s MineThatData Email data [hillstrom2008]. The Hillstrom dataset contains records of 64,000 customers who purchased within the last twelve months. They were part of an e-mail campaign to assess the effectiveness of distinct campaign strategies. Two treatment groups, “Men’s” and “Women’s” email campaigns, and a control group were established. Treatments were randomly assigned. Our analysis primarily focuses on a combined treatment group, which constitutes roughly 66% of the dataset. While the original dataset has different outcomes, we looked at the dollars spent in the two weeks post-campaign. The dataset provides data on recent purchase patterns (Recency), annual spending categories (History Segment) and values (History), merchandise type, either Mens (Mens) or Womens (Womens), geographical location via zip code (Zip Code), newcomer status (Newbie), and purchasing avenues (Channel). After subsampling, we end up with a randomized trial of size $n_{\operatorname{rct}}=7680$ and an observational dataset of size $n_{\operatorname{os}}=20500$. 2.3 is enforced by excluding urban zip codes in the trial.
@@
  • VOTE dataset [gerber2008social]. The VOTE dataset studies the effect of social pressure on voting behaviors among Michigan’s registered voters, focusing on those who voted in the prior election and met certain criteria. The primary outcome is a binary variable indicating whether the letter recipients voted. In this randomized trial, participants were allocated to a control group or one of four treatment groups. The treatment groups received distinct letters, each varying in social pressure intensity, aimed at encouraging voting. The most persuasive letter provided insight into the recipient’s neighbors’ voting patterns from the previous two elections and implied updates on neighbors’ subsequent voting actions in future letters. Using the split in [stutz2023], we incorporated roughly 190,000 samples in the control group, and we kept the treatment group with the strongest letter, leaving about 38,000 samples. We retained preprocessed features like age, household size, gender, and two scores reflecting past voting habits and the voting patterns of neighbors. After subsampling, we end up with a randomized trial of size $n_{\operatorname{rct}}=10650$ and an observational dataset of size $n_{\operatorname{os}}=36800$. We discard households with more than 4 participants to enforce 2.3.
@@
  • Tennessee STAR Project [word1990state]. The Tennessee STAR experiment, initiated in 1985, was a randomized study examining the impact of class size on students’ standardized test scores, tracking them from kindergarten through third grade. Initially, students and teachers were randomly placed into class sizes, intending to maintain these conditions throughout the study. We follow the dataset preprocessing outlined in [kallus2018removing]. Their analysis concentrates on two conditions: small classes (13-17 students) and regular-sized classes (22-25 students). They used the class size at first grade as the treatment variable, observing 4,509 students. Their outcome aggregates scores from listening, reading, and math tests at the end of the first grade. After excluding students with missing values, the final sample consisted of 4,218 students: 1,805 in small classes (treatment) and 2,413 in regular-sized classes (control). The observed features for each student are gender, race, birth month, birthday, birth year, free lunch given or not and teacher ID. After subsampling, we end up with a randomized trial of size $n_{\operatorname{rct}}=600$ and an observational dataset of size $n_{\operatorname{os}}=1800$. For the STAR project, we keep inner-city and suburban students but remove urban and rural ones to enforce 2.3.
@@
We use QB [dorn2022sharp] for the test $\hat{\phi}_{\widetilde{\mathrm{os}}}$. For continuous outcomes (Hillstrom and STAR), we fit a random forest regressor for the quantile functions, while we leverage the closed-form solution for the quantiles in the binary case (VOTE). For the test $\hat{\phi}_{\mathrm{rct}}$ we use the B-Learner [oprescu2023b], fitting respective random forest regressors for the quantiles, the outcome and the bounds models. For the binary case, we use the closed-form solution for the quantiles and fit the outcome and bounds models with XGBoost [chen2015xgboost]. We always train a logistic regressor for the propensity score. We report mean and standard error over 15 runs and set $n_{\text{bootstrap}}=200,\alpha=0.05$ for all experiments.
@@ B.3 Women’s Health Initiative
We binarize a composite outcome, called the “global index”, in our analysis, where Y = 1 if coronary heart disease or stroke was observed in the first seven years of follow-up, and Y = 0 otherwise. Note that Y = 0 could also occur from censoring. To establish treatment and control groups in the observational study, we use questionnaire data in which participants confirm or deny usage of combination hormones (i.e. both estrogen and progesterone) in the first three years. Using this procedure, we end up with a total of $n_{\operatorname{os}}=33,511$ patients. Finally, we restrict the set of covariates used to those that are measured in both the RCT and the observational study. In particular, we use as covariates only those measured in both the RCT and observational study, and we further restrict them to those identified as significant in epidemiological literature, such as in [prentice2005combined]. Specifically, the covariates in our analysis are: AGE, ETHNIC_White, BMI, SMOKING_Past_Smoker, SMOKING_Current_Smoker, EDUC_x_College_graduate_or_Baccalaureate Degree, EDUC_x_Some_post-graduate_or_professional, MENO, PHYSFUN. The data used is available on BIOLINCC.
@@
We train logistic regression for both outcome models and propensity score. We use as sensitivity bounds DVDS [dorn2021doubly] for the test $\hat{\phi}_{\widetilde{\mathrm{os}}}$, and B-Learner for the test $\hat{\phi}_{\mathrm{rct}}$. We test for confounding in one direction, i.e. we only compute the test statistic $\hat{T}_{\Gamma}^{+}$. We set $n_{\text{bootstrap}}=100$ and $\alpha=0.05$ for all experiments.